OpenStudy (anonymous):

At equilibrium, a 1.00 M OCl− solution has an [OH−] of 5.75 × 10−4 M. Which of the following is the correct pH of the solution? 3.24 3.3 × 10−7 14.00 10.76

OpenStudy (rushwr):

\[[OH ^{-}] = 5.75 * 10^{-4}\] we know pKw = 10 ^-14 \[[H ]= \frac{ Kw }{ [OH] } = \frac{ 10^{-14} }{ 5.75 * 10^{-4} }= 0.174 * 10^{-10}\]\[pH = -\log _{10} [H ^{+}]\] pH = 10.76